OpenStudy (anonymous):

What does this question mean, someone help please! The question is: Solve: 24 -one over fourN = -4, please help! Thanks, i'll give a medal for help! :D.

OpenStudy (anonymous):

1 4 for one over four and then it shows the variable N :l.

OpenStudy (anonymous):

\[\frac{4-1}{N}=-4\]

OpenStudy (anonymous):

What does N equal? :l.

OpenStudy (anonymous):

u need help solve for N?

OpenStudy (anonymous):

I'm not too good at this stuff >.< :s..

OpenStudy (anonymous):

N*-4=4-1 -4N = 3 N = 3/-4

OpenStudy (anonymous):

\[N=\frac{-3}{4}\]

OpenStudy (anonymous):

wait i wrong

OpenStudy (anonymous):

\[\frac{24-1}{4N}=-4\] i missing

OpenStudy (anonymous):

24 - 1 ------ = -4 isnt it like this? 4N 23 ----=N -16 and therefore N= -23/16 ?

OpenStudy (anonymous):

-4*4N=24-1 -16N =23 \[N=\frac{-16}{23}\]

OpenStudy (anonymous):

I missind 24 and fourN

OpenStudy (anonymous):

ok so i wAS right.... lol

OpenStudy (anonymous):

Thanks, medal for you ^_^.

OpenStudy (anonymous):

ok thnx... Nancy was right also in her solution except she missed the 24

OpenStudy (anonymous):

good luck skiller

OpenStudy (anonymous):

the answer will get exactly if u use caculator pi instate use 3.14

OpenStudy (anonymous):

Another interpretation of the problem statement: \[24-\frac{1}{4n}=-4 \]In this case\[n=\frac{1}{112} \]